OpenStudy (anonymous):

y=e^(sqrtx). I need to find the inner and outer function and the derivative of that. calculus problem. I need help with exponents guys.

OpenStudy (anonymous):

is e^x outer and sqrtx the inner?

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

ok what next?

OpenStudy (blurbendy):

use the chain rule on e^(sqrt[x]) let u = sqrt[x] \[\frac{ d }{ du }e^u = e^u\] now \[e^{\sqrt{x}}\frac{ d }{ du }\sqrt{x}\] Use the power rule (sqrt[x] = x^(1/2) \[\frac{ d }{ du }x^{1/2} = \frac{ x^{-1/2} }{ 2 }\] = \[\frac{ 1 }{ 2\sqrt{x} }e^{x}\]

OpenStudy (anonymous):

thanks!

OpenStudy (blurbendy):

np